OpenStudy (anonymous):

An isosceles triangle has base B and height H. Assuming its area remains constant, express the rate of change of the base with respect to the height.

OpenStudy (anonymous):

A= 1/2 b * h 2A = b * h b = 2A/h db/dh 2A/h = db/dh 2A * h^-1 = 2A * -h^-2 = -2A/h^2